### 22 Apr 2026 v1

# Cesàro-Hardy operators on $L^{p}[0,1]$: fine spectrum, weighted Koopman semigroups and invariant subspaces

Abstract.
In this paper we study boundedness and detailed spectral properties for the Cesàro-Hardy operator and some generalizations in $L^{p}[0,1]$. The study employs $C_{0}$-semigroup theory, expressing the Cesàro-Hardy operators and their dual operators through subordination with $C_{0}$-semigroups $T(t)$ and $S(t)$ respectively. The spectral properties of the semigroup’s infinitesimal generators are transferred to the Cesàro-Hardy operators using functional calculus methods. Furthermore, some implications for the Invariant Subspace Problem are explored by demonstrating the universality of certain translations related to the semigroup $T(t),$ and providing results on the invariant subspaces of these operators.
Key words and phrases:
Cesàro-Hardy operators, fine spectrum, weighted Koopman semigroups, Invariant Subspace Problem, universal operators
2020 Mathematics Subject Classification:
47A10, 47A15, 47B33, 47D03, 47G10
Authors have been partially supported by Grant PID2022-137294NBI00 funded by MICIU/AEI/10.13039/501100011033 and “ERDF/EU” and Project E48-20R, Gobierno de Aragón, Spain. The second author has been partially supported by Grant PRE-2020-092311 funded by MICIU/AEI/10.13039/501100011033 and, by “ESF Investing in your future”.
1. Introduction
Our starting point is the work of A. Brown, P. R. Halmos and A. L. Shields, [16]. In their paper, the Cesàro-Hardy operators, $C_{0},C_{1}$ and $C_{\infty}$ were defined on $\ell^{2},L^{2}[0,1]$ and $L^{2}(0,\infty)$ respectively by
$C_{0}f(n)=\frac{1}{n+1}\sum_{j=0}^{n}f(j),\quad C_{1}f(x)=\frac{1}{x}\int_{0}^{x}f(s)\,ds\quad\text{and}\quad C_{\infty}f(x)=\frac{1}{x}\int_{0}^{x}f(s)\,ds.$
These operators are bounded due to the Hardy inequality (see [28, Section 9.8]). Furthermore, the authors characterized their spectrum and point spectrum, offering insights beyond simple boundedness. Building on this foundation, D. W. Boyd [15] extended these spectral results for the operator $C_{\infty}$ to the spaces $L^{p}(0,\infty)$ for $1<p\leq\infty$, while G. Leibowitz followed suit with the operators $C_{0}$ and $C_{1}$ on the spaces $\ell^{p}$ and $L^{p}[0,1]$ for $1<p\leq\infty$ respectively [29, 30]. In [34], the author computed the approximate and continuous spectrum of $C_{1},$ and in [24] the essential spectrum of $C_{0},C_{1}$ and $C_{\infty}$ was studied.
Since then, the spectrum and point spectrum of the Cesàro-Hardy operators have been extensively studied across various function spaces, including $C[0,1]$, $C_{l}[0,\infty]$, $C(0,\infty)$, $L^{p}_{\text{loc}}(0,\infty)$ with $1<p<\infty$ [6], $C^{\infty}(0,\infty)$ [8, 9], and ultra-differentiable function spaces [11]. Recent studies have also examined $C_{1}$ in Lorentz spaces $L_{p,q}(0,1)$ for $1<p<\infty$ and $1\leq q\leq\infty$ [45] and in rearrangement invariant spaces over a finite interval and a half line [5].
The discrete Cesàro-Hardy operator’s spectrum and fine spectra have similarly been analyzed in various spaces, including $\ell_{p}$ $(1<p<\infty)$ [26, 19], $c_{0}$ [38, 4], $\ell_{\infty}$, Bachelis spaces $N_{p}$ $(1<p<\infty)$ [21], $bv_{0}$, and bvbv [35, 36], as well as in weighted $\ell_{p}$ spaces [10, 7], discrete Cesàro spaces $\operatorname{ces}_{p}$ $(p\in\{0\}\cup(1,\infty))$ [20], and their dual spaces $d_{s}$ $(1<s<\infty)$ [14].
Subsequently, in 2014, C. Lizama, P. J. Miana, R. Ponce and L. Sánchez-Lajusticia studied in [31] a generalized Cesàro-Hardy operator. For $\alpha>0$, they considered the Cesàro-Hardy operator of order $\alpha$ defined by
$\mathcal{C}_{\alpha}f(s)=\frac{\alpha}{s^{\alpha}}\int_{0}^{s}(s-u)^{\alpha-1}f(u)\,du,\quad s\in(0,\infty),$
on certain weighted Sobolev spaces $\mathscr{T}_{p}^{\beta}(t^{\beta})$, where $p\geq 1$ and $\beta>0$. The space $\mathscr{T}_{p}^{\beta}(t^{\beta})$ consists of functions on the positive half-line $(0,\infty)$ such that $t\to t^{\beta}W^{\beta}_{+}f(t)$, where $W^{\beta}_{+}f$ is the Weyl fractional derivative of $f$ of order $\beta$, belongs to $L^{p}(0,\infty)$. Note that these spaces are contained within the Lebesgue spaces $L^{p}(0,\infty)$ (they coincide for $\beta=0$). The significance of this work lies not in the generalization of the operator itself, but in the approach employed. The introduction of $C_{0}$-semigroup theory to this problem offers a clearer understanding. Specifically, the approach involved representing the generalized Cesàro-Hardy operator as an operator subordinated via a $C_{0}$-group, allowing the spectral analysis of the (semi)group’s generator to be applied directly to the operator itself. This work can be viewed as a natural continuation of Boyd’s work [15]. Later, the discrete case was revisited within the $C_{0}$-semigroup framework in [1] extending Leibowitz’s work [29]. Therefore, the only case left was the study of this generalized Cesàro-Hardy operator on the unit interval using $C_{0}$-semigroup techniques, which is where this work starts.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$, consider the Cesàro-Hardy operator of order $\alpha$ given by
$\mathcal{C}_{\alpha}f(s)=\frac{\alpha}{s^{\alpha}}\int_{0}^{s}(s-u)^{\alpha-1}f(u)\,du,\quad s\in(0,1),$
and the dual Cesàro-Hardy operator of order $\alpha$ given by
$\mathcal{C}_{\alpha}^{*}f(s)=\alpha\int_{s}^{1}\frac{(u-s)^{\alpha-1}}{u^{\alpha}}f(u)\,du,\quad s\in(0,1),$
which will be key to prove some spectral properties. Note that $\mathcal{C}_{1}=C_{1}.$ For $\mathcal{C}_{\alpha},$ the change of variables $u=e^{-t}s$ yields
$\mathcal{C}_{\alpha}f(s)=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}e^{-t}f(e^{-t}s)\,dt,$
and for $\mathcal{C}_{\alpha}^{*}$ the change of variable $u=e^{t}s$ gives
$\mathcal{C}_{\alpha}^{*}f(s)=\alpha\int_{0}^{\log(1/s)}(1-e^{-t})^{\alpha-1}f(e^{t}s)\,dt=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}\chi_{(0,e^{-t})}(s)f(e^{t}s)\,dt,$
where $\chi_{I}(s)$ denotes the characteristic function of the interval II. The operator families, $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ given by
(1.1) $T(t)f(s):=e^{-t}f(e^{-t}s),\quad S(t)f(s):=\chi_{[0,e^{-t})}(s)f(e^{t}s),\quad s\in[0,1],\,t\geq 0,$
are $C_{0}$-semigroups on $L^{p}[0,1]$ for $1\leq p<\infty$, see Theorem $2.$ Both semigroups are referred to in the literature as Koopman or composition semigroups. Next, with this representation in mind, we obtain the fine spectrum of both infinitesimal generators and transfer their spectral properties, first to the semigroup, using standard $C_{0}$-semigroup theory, see Theorems 4.1 and 4.2. And later on, to the Cesàro-Hardy operator using the Hille-Phillips functional calculus and the functional calculus for sectorial operators developed by M. Haase in [27], see Theorems 5.3 and 5.4. This semigroup approach allows us to get, in a unified way, the spectrum of wide class of integral operators, see Section 5.
As a consequence of the spectral properties we prove in Theorem 7.4 that for $t>0$ and $\mu\in B(0,e^{-t/2})$, the open ball centered at 0 with radius $e^{-t/2}$, the operator $\mu-T(t)$ is universal on $L^{2}[0,1]$ in the sense of Rota, see [39, 40]. This notion of universality is intimately connected to the well-known Invariant Subspace Problem, a relationship that we will clarify in the following lines. In an infinite-dimensional separable complex Hilbert spaces $H$, a bounded linear operator $U$ is universal if for every operator $T$ in $\mathcal{B}(H)$ (the algebra of bounded linear operators acting on $H$) there exist a scalar $\lambda$, a closed invariant subspace $M$ of $U$, and an isomorphism $\Theta:H\to M$ such that $\Theta T=\lambda U\mid_{M}\Theta$. Essentially, $T$ is similar to a scaled restriction of $U$ to one of its invariant subspaces. For a universal operator $U$, the following statements are equivalent:
  • (i)
    every bounded linear operator on an infinite-dimensional separable Hilbert space has a non-trivial closed invariant subspace.
  • (ii)
    every minimal non-trivial invariant subspace of $U$ is one-dimensional.
S. R. Caradus [17] showed that if an operator $T$ in a separable Hilbert space has an infinite-dimensional kernel and is surjective, then $T$ is universal. We apply this result to prove the universality of the translations of $T(t)$ in Theorem 7.4.
Recently, J. Agler and J. E. McCarthy have characterized the invariant subspaces of $\mathcal{C}_{1}$ on $L^{2}[0,1].$ They prove that the lattice of non-zero invariant subspaces of $\mathcal{C}_{1}$ on $L^{2}[0,1]$ (denoted by $\operatorname{Lat}\mathcal{C}_{1}$), is the set of Monomial spaces (see more details about this concept in Section 7). We prove in Theorem 7.6 that the lattice of the closed invariant subspaces of the semigroup $(T(t))_{t>0}$ coincides with the one of $\mathcal{C}_{1}$ on $L^{2}[0,1],$ that is, $\cap_{t>0}\operatorname{Lat}T(t)=\operatorname{Lat}\mathcal{C}_{1}.$
While we initially presented the work of [31] as the motivation for addressing the problem discussed in this paper, it is important to acknowledge that the earliest known reference establishing the connection between $C_{0}$-semigroups and the Cesàro-Hardy operator is the article by $C.$ Cowen [18]. Afterward, the technique was further employed by A. Arvanitidis and A. Siskakis [13], who proved that the half-plane versions of Cesàro-Hardy operators on Hardy spaces are bounded. Similarly, W. Arendt and B. de Pagter [12] examined the Cesàro-Hardy operator defined in an interpolation space $E$ between $L^{1}$ and $L^{\infty}$ on the positive real half-line. Additionally, as we mentioned before, the generalized discrete Cesàro-Hardy operator was studied in [1], and more recently in [2] the authors used the $C_{0}$-semigroup approach to determine the fine spectrum of generalized Hausdorff operators on various analytic spaces, including a generalized Cesàro-Hardy operator.
The paper is structured as follows. In Section 2, we examine the properties of the one-parameter families $T(t)$ and $S(t)$. Section 3 discusses the spectral properties of their infinitesimal generators. This sets the stage for a detailed analysis of the spectrum of the semigroups and the generalized Cesàro-Hardy operator in Section 4 and Section 5. In section 6, we present a nice connection between the $C_{0}$-semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ on $L^{p}[0,1]$ and the $C_{0}$-semigroups on $\ell^{p}$ introduced in [1] to study the discrete Cesàro operators. These connections allow us to show new intertwining properties between Cesàro-Hardy operators $\mathcal{C}_{\alpha},\mathcal{C}_{\alpha}^{*}$ and the discrete generalized Cesàro operators on $\ell^{p}$. Finally, Section 7 addresses the universality of the translations of $T(t)$ and the Invariant Subspace Problem. Moreover, in this last section we give some results about the invariant subspaces of the semigroup $(T(t))_{t>0}$, and also about the cyclic subspaces of every universal operator $T(t).$
To conclude this introduction, we recall some classical definitions. Let $X$ be a Banach space, $\mathcal{B}(X)$ will denote the Banach algebra of bounded linear operators on $X$ with the operator norm $\left\|\cdot\right\|$. Let $1\leq p\leq\infty$, the Lebesgue spaces $L^{p}[0,1]$ are Banach spaces with the $L^{p}$-norm denoted by $\left\|\cdot\right\|_{p}$. For $1\leq p<\infty$ the dual space of $L^{p}$ is $L^{p^{\prime}}$, where $p^{\prime}$ denotes the conjugate exponent of pp. For $1<p<\infty$, $p^{\prime}$ satisfies $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. If $p=1$, then $p^{\prime}=\infty$ and $\frac{1}{p^{\prime}}$ is understood as 0. Throughout the paper we will maintain this notation.
Lastly, we recall some classical definitions in spectral theory. For a closed linear operator $\mathcal{A}:D(\mathcal{A})\subseteq X\to X$ on a Banach space $X$, where $D(\mathcal{A})$ denotes the domain of the operator, we denote the spectrum by $\sigma(\mathcal{A})$, the point spectrum by $\sigma_{point}(\mathcal{A})$, the approximate point spectrum by
[…]
the residual spectrum by
$\sigma_{r}(\mathcal{A}):=\{\lambda\in\mathbb{C}\,:\,\operatorname{Ran}(\lambda-\mathcal{A})\text{ is not dense in }X\},$
and the essential spectrum by
$\sigma_{ess}(\mathcal{A}):=\{\lambda\in\mathbb{C}\,:\,\dim(\ker(\lambda-\mathcal{A}))=\infty\text{ or }\operatorname{codim}(\operatorname{Ran}(\lambda-\mathcal{A}))=\infty\}$
where $\operatorname{codim}(Y):=\dim(X/Y)$ for any subspace $Y\subseteq X$. The spectral radius $r(\mathcal{A})$ of a bounded operator $\mathcal{A}$ is given by $r(\mathcal{A}):=\sup\{\left|\lambda\right|\,:\,\lambda\in\sigma(\mathcal{A})\}$.
2. Weighted Koopman semigroups on $L^{p}[0,1]$
In this section we study, for $1\leq p\leq\infty,$ the $C_{0}$-semigroup structure of the two operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ on the spaces $L^{p}[0,1]$ defined on (1.1). Along with this study, several results about the integral operators, $\Lambda_{0}^{\lambda},\Lambda_{1}^{\lambda}$ and $L^{\lambda}$ defined on (2.1) and (2.2) are stated.
Proposition 2.1.
Let $1\leq p\leq\infty$ and $t\geq 0$. Then:
  1. (i)
    The operator $T(t)$ is bounded on $L^{p}[0,1]$ and $\left\|T(t)\right\|=e^{-t/p^{\prime}}$.
  2. (ii)
    The operator $S(t)$ is bounded on $L^{p}[0,1]$ and $\left\|S(t)\right\|=e^{-t/p}$.
Proof.
For $1\leq p<\infty$ the operator $T(t)$ satisfies
$\left\|T(t)f\right\|_{p}^{p}=e^{-pt}\int_{0}^{1}\left|f(e^{-t}s)\right|^{p}\,ds=e^{(1-p)t}\int_{0}^{e^{-t}}\left|f(u)\right|\,du\leq e^{(1-p)t}\left\|f\right\|_{p},$
and if $p=\infty$
$\left\|T(t)f\right\|_{\infty}=\operatorname*{ess\,sup}_{s\in[0,1]}\left|e^{-t}f(e^{-t}s)\right|\leq e^{-t}\left\|f\right\|_{\infty},$
so $\left\|T(t)\right\|\leq e^{-t/p^{\prime}}$ for $t\geq 0$. Moreover, $\left\|T(t)\right\|=e^{-t/p^{\prime}}$ since the equality holds for the characteristic function of the interval $0,e^{-t})$ for $t\geq 0$. If $1\leq p<\infty$ the operator $S(t)$ satisfies
$\left\|S(t)f\right\|_{p}^{p}=\int_{0}^{e^{-t}}\left|f(e^{t}s)\right|^{p}\,ds=e^{-t}\int_{0}^{1}\left|f(u)\right|^{p}\,du=e^{-t}\left\|f\right\|_{p},$
and if $p=\infty$
$\left\|S(t)f\right\|_{\infty}=\operatorname*{ess\,sup}_{s\in[0,e^{-t}]}\left|f(e^{t}s)\right|=\left\|f\right\|_{\infty},$
so $\left\|S(t)\right\|=e^{-t/p}$ with $t\geq 0$. ∎
Theorem $2.$
Let $1\leq p<\infty$. The one-parameter operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ are $C_{0}$-semigroups on $L^{p}[0,1]$ with growth bounds $1/p^{\prime}$ and 1/p-1/p, respectively. Their infinitesimal generators, AA and $B$, are given by
$Af(s)=-sf^{\prime}(s)-f(s),\quad\text{ and }\quad Bf(s)=sf^{\prime}(s),\quad s\in[0,1],$
defined on the domains,
$D(A):=\{f\in L^{p}[0,1],f\in AC_{loc}(0,1)\text{ and }sf^{\prime}(s)\in L^{p}[0,1]\},$
and
$D(B):=\{f\in L^{p}[0,1],f\in AC_{loc}(0,1)\text{, }\lim_{s\to 1^{-}}f(s)=0\text{ and }sf^{\prime}(s)\in L^{p}[0,1]\},$
where $AC_{loc}(0,1)$ is the space of locally absolutely continuous functions on $(0,1)$.
Proof.
Firstly, note that both semigroups satisfy the semigroup property and are bounded operators for each $t\geq 0$ as shown in Proposition 2.1. For strong continuity, this condition is fulfilled due to the dominated convergence theorem.
Next, the properties stated for the generators are proven. Notice that for the semigroup $T(t)$ we can compute the generator directly obtaining $Af(s)=-sf^{\prime}(s)-f(s)$ for a.e. $s\in[0,1]$ with domain
$D(A):=\{f\in L^{p}[0,1],f\in AC_{loc}(0,1)\text{ and }sf^{\prime}(s)\in L^{p}[0,1]\}.$
[…]
Now we define the space $Y$ as
$Y=\{f\in L^{p}(0,\infty):f(x)=0\text{ a.e. for }x>1\}.$
[…]
$SS(t)f(x)=\begin{cases}f(xe^{t}),&0\leq xe^{t}\leq 1\\ 0,&xe^{t}>1\end{cases}=\chi_{[0,e^{-}t]}(x)f(xe^{t})$
[…]
$D(\mathcal{B}_{Y})$ $=\{f\in D(\mathcal{B})\cap Y:\mathcal{B}f\in Y\}$
$=\{f\in Y\,:\,f\in AC_{loc}(0,\infty)\text{ and }xf^{\prime}(x)\in Y\}$
$=\{f\in Y\,:\,f\in AC_{loc}(0,1)\text{, }\lim_{x\to 1}f(x)=0\text{ and }xf^{\prime}(x)\in Y\},$
where the last equality follows from the definition of local absolute continuity and the definition of $Y$.
[…]
$Bf(s):=(J^{-1}\mathcal{B}_{Y}J)(s)=sf^{\prime}(s)\quad\text{a.e. }s\in[0,1]$
with domain
$D(B)$ $=\{f\in L^{p}[0,1]:Jf\in D(\mathcal{B}_{Y})\}$
$=\{f\in L^{p}[0,1]:f\in AC_{\text{loc}}(0,1),\lim_{s\to 1^{-}}f(s)=0,\text{ and }sf^{\prime}(s)\in L^{p}[0,1]\}.$
as we wanted to prove. ∎
Remark 2.3.
Observe that $S(t)$ on $L^{p^{\prime}}[0,1]$ is the dual operator of $T(t)$ on $L^{p}[0,1]$ for $1\leq p<\infty$,
$\int_{0}^{1}T(t)f(s)g(s)\,ds$ $=\int_{0}^{1}e^{-t}f(e^{-t}s)g(s)\,ds=\int_{0}^{1}f(u)\chi_{[0,e^{-t})}(u)g(e^{t}u)\,du$
$=\int_{0}^{1}f(u)(S(t)g)(u)\,du.$
Remark 2.4.
Notice that the semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ are not strongly continuous on $L^{\infty}[0,1]$ as they are not uniformly continuous, see [32, Theorem 4].
In the following, we introduce three integral operators which play a key role in the study of the spectral properties. Let $\lambda\in\mathbb{C}$ and $f$ a measurable function on $[0,1],$ we write
(2.1) $(\Lambda_{0}^{\lambda}f)(s)=\frac{1}{s^{\lambda+1}}\int_{0}^{s}u^{\lambda}f(u)\,du,\quad(\Lambda_{1}^{\lambda}f)(s)=s^{\lambda}\int_{s}^{1}\frac{f(u)}{u^{\lambda+1}}\,du,\quad s\in(0,1),$
and
(2.2) $L^{\lambda}f=\int_{0}^{1}\frac{f(u)}{u^{\lambda+1}}\,du.$
Proposition 2.5.
Let $1\leq p<\infty$ and $f\in L^{p}[0,1],$ then
  1. (i)
    $\Lambda_{0}^{\lambda}f=\int_{0}^{\infty}e^{-\lambda t}T(t)f\,dt$ in $L^{p}[0,1]$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime}.$
  2. (ii)
    $\Lambda_{1}^{\lambda}f=\int_{0}^{\infty}e^{-\lambda t}S(t)f\,dt$ in $L^{p}[0,1]$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p$.
Proof.
Let $f\in L^{p}[0,1]$, by Hölder’s inequality one gets
$\int_{0}^{s}|u^{\lambda}f(u)|\,du\leq\|f\|_{p}\biggl(\int_{0}^{s}u^{p^{\prime}{\mathfrak{R}}{\mathfrak{e}}\,\lambda}\,du\biggr)^{1/p^{\prime}},\quad s\in(0,1),$
which converges for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime},$ and then $\Lambda_{0}^{\lambda}f$ is a measurable function. Also,
$\int_{s}^{1}\left|\frac{f(u)}{u^{\lambda+1}}\right|\,du\leq\|f\|_{p}\biggl(\int_{s}^{1}\frac{1}{u^{p^{\prime}({\mathfrak{R}}{\mathfrak{e}}\,\lambda+1)}}\,du\biggr)^{1/p^{\prime}},\quad s\in(0,1),$
which always converges, so $\Lambda_{1}^{\lambda}f$ is a measurable function. Notice that for the case $p=1$ the last integral on each inequality has to be understood as the $L^{\infty}$ norm.
By the growth bounds of the semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ given in Theorem 2.2, the right-side hand integrals in the statements of this proposition are Bochner convergent in $L^{p}[0,1].$
Finally, for $s\in(0,1),$ the change of variable $u=e^{-t}s$ gives
$(\Lambda_{0}^{\lambda}f)(s)=\int_{0}^{\infty}e^{-\lambda t}(T(t)f)(s)\,dt,$
and the change of variable $u=e^{t}s$ implies
$(\Lambda_{1}^{\lambda}f)(s)=\int_{0}^{\infty}e^{-\lambda t}(S(t)f)(s)\,dt.\qed$
Now, we turn our attention to the functional $L^{\lambda}.$ Depending on the values of $p$ and $\lambda$ the domain will change and also the behavior of our functional. This is shown in the following results.
Lemma 2.6.
Let $1\leq p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p.$ Then $L^{\lambda}:L^{p}[0,1]\to\mathbb{C}$ is a continuous linear functional on $L^{p}[0,1].$ Additionally, for $p=1$ and ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1$, $L^{\lambda}$ is also continuous.
Proof.
Let $f\in L^{p}[0,1]$ and $\lambda\in\mathbb{C}$ such that ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p.$ Then by Hölder’s inequality one gets
$\left|\int_{0}^{1}\frac{f(u)}{u^{\lambda+1}}\,du\right|\leq\left\|f\right\|_{p}\left(\int_{0}^{1}\frac{1}{u^{p^{\prime}({\mathfrak{R}}{\mathfrak{e}}\,\lambda+1)}}\right)^{1/p^{\prime}},$
which is convergent, therefore $L^{\lambda}$ is a well-defined continuous linear functional on $L^{p}[0,1].$ Notice that for the case $p=1$ the last integral has to be understood as the $L^{\infty}$ norm and therefore for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1$ we also have that $L^{\lambda}$ is a well-defined continuous linear functional on $L^{1}[0,1].$
The last result leaves unexplored the case when ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p$. In particular, notice that $L^{\lambda}$ might not even be well-defined when $p\neq 1$. Nonetheless, we can define the functional on a dense subspace of $L^{p}[0,1]$, $C_{c}^{\infty}(0,1)$ the space of infinite differentiable functions of compact support on $(0,1)$. In that case, see that the functional is unbounded.
Lemma 2.7.
Let $1<p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p.$ Then $L^{\lambda}:C_{c}^{\infty}(0,1)\to\mathbb{C}$ is an unbounded linear functional with the norm $\|\cdot\|_{p}.$
Proof.
First, notice that $L^{\lambda}$ is well-defined for all the functions on $C_{c}^{\infty}(0,1)$ as they have compact support. To check that it’s unbounded we define the sequence
$f_{n}(u)=\chi_{[\frac{1}{n},\frac{1}{2}]}(u)\frac{u^{\lambda}}{\log(1/u)^{\frac{1}{p}}},\quad u\in[0,1],\;n>2.$
On one hand, we have that,
$L^{\lambda}f_{n}=\int_{\frac{1}{n}}^{\frac{1}{2}}\frac{du}{u\log(1/u)^{\frac{1}{p}}}=\int_{\log 2}^{\log n}s^{-\frac{1}{p}}\,ds=p^{\prime}\left((\log n)^{\frac{1}{p^{\prime}}}-(\log 2)^{\frac{1}{p^{\prime}}}\right).$
On the other hand,
$\left\|f_{n}\right\|_{p}^{p}=\int_{\frac{1}{n}}^{\frac{1}{2}}\left|\frac{u^{\lambda}}{\log(1/u)^{\frac{1}{p}}}\right|^{p}\,du=\int_{\frac{1}{n}}^{\frac{1}{2}}\frac{du}{u\log(1/u)}=\int_{\log 2}^{\log n}\frac{1}{s}\,ds=\log(\log n)-\log(\log 2).$
Therefore,
$\lim_{n\to\infty}\frac{\left|L^{\lambda}f_{n}\right|}{\left\|f_{n}\right\|_{p}}=\lim_{n\to\infty}\frac{p^{\prime}(\log n)^{\frac{1}{p^{\prime}}}}{\log(\log n)^{\frac{1}{p}}}=+\infty,$
So, the functional $L^{\lambda}$ is unbounded. ∎
Remark 2.8.
Note that for $1\leq p<\infty$ and ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p,$ the kernel of the functional $L^{\lambda}$ on $C_{c}^{\infty}(0,1)$ is not the whole domain as it’s not the zero functional.
Corollary 2.9.
For $1<p<\infty$ and ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p,$ the kernel of the functional $L^{\lambda}$ on $C_{c}^{\infty}(0,1)$ is dense on $L^{p}[0,1].$
Proof.
Recall that a non-zero linear functional on a topological vector space $X$, which is the case for $L_{\lambda}$ on $C_{c}^{\infty}(0,1)$ with the norm $\|\cdot\|_{p}$, is continuous if and only if the kernel is not dense in $X$, see [41, Theorem 1.18]. So, the result follows by Lemma 2.7 and the density of $C_{c}^{\infty}(0,1)$ in $L^{p}[0,1]$ with the norm $\|\cdot\|_{p}.$
3. Spectra of differential operators
In this section we present the fine spectrum of the infinitesimal generators AA and $B$ on $L^{p}[0,1].$
Proposition 3.1.
Let $1\leq p<\infty.$ The point spectrum of $B$ is empty, that is, $\sigma_{point}(B)=\emptyset.$ The point spectrum of AA is the set $\sigma_{point}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}\}.$ Furthermore, the eigenspace of each $\lambda\in\sigma_{point}(A)$ is one-dimensional and generated by $g_{\lambda}(s):=\frac{1}{s^{\lambda+1}}.$
Proof.
Let $\lambda\in\mathbb{C}$ and $f\in D(B)$ such that $Bf=\lambda f$. Then, $f$ is a solution of the differential equation
$sf^{\prime}(s)-\lambda f(s)=0.$
The nonzero solutions to this equation have the form $f(s)=cs^{\lambda}$ with $c\in\mathbb{C}$ and $c\neq 0,$ but these functions are not in $D(B)$ because $\lim_{s\to 1^{-}}f(s)\neq 0.$ Therefore, $\sigma_{point}(B)=\emptyset.$
If $\lambda\in\mathbb{C}$ and $f\in D(A)$ such that $Af=\lambda f$, then $f$ satisfies
$(\lambda+1)f(s)+sf^{\prime}(s)=0.$
The solutions are of the form $f(s)=cg_{\lambda}(s)$, where $c$ is a complex constant. These solutions belong to $L^{p}[0,1]$ if and only if $c=0$ or ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}$. Therefore, $\sigma_{point}(A)=\{\lambda\in\mathbb{C}:{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}\}$, and the associated eigenspace is the one-dimensional subspace generated by $g_{\lambda}$. ∎
Proposition 3.2.
Let $1\leq p<\infty.$ The spectrum of AA is the set $\sigma(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}.$ Moreover, for $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime},$
$(\lambda-A)^{-1}f=\Lambda_{0}^{\lambda}f,\quad f\in L^{p}[0,1].$
Proof.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p^{\prime}$, by the integral representation of the resolvent operator of AA in terms of the semigroup $(T(t))_{t\geq 0}$ (see e.g. [23, Th. II.1.10]) and Proposition 2.5, one has
$(\lambda-A)^{-1}f=\Lambda_{0}^{\lambda}f,\quad f\in L^{p}[0,1].$
Therefore, $\sigma(A)\subseteq\{\lambda\in\mathbb{C}:{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}$. Since $\sigma_{point}(A)\subseteq\sigma(A)$ and the spectrum is a closed subset of $\mathbb{C},$ the result follows by Proposition 3.1. ∎
Proposition $3.$
Let $1\leq p<\infty.$ The spectrum of $B$ satisfies $\sigma(B)\subseteq\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p\}.$ Moreover, for $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda>-1/p,$
$(\lambda-B)^{-1}f=\Lambda_{1}^{\lambda}f,\quad f\in L^{p}[0,1].$
Proof.
The result is a straightforward consequence of [23, Th. II.1.10] and Proposition 2.5. ∎
Remark 3.4.
As we have seen in Remark 2.3 the semigroup $S(t)$ is the dual semigroup of $T(t)$ for each $t\geq 0$. Notice that as the space $L^{p}[0,1]$ is reflexive for $1<p<\infty$ using $C_{0}$-semigroup theory and duality we could compute the exact spectrum directly of the operator $B$, see [23, Paragraph II.2.5]. But, in order to do a detailed study we leave this as a remark.
We proceed to study the range space of the operator $\lambda-A:D(A)\to L^{p}[0,1]$ for a fixed $\lambda\in\mathbb{C}$. Given $f\in L^{p}[0,1],$ it is an easy computation that the solutions of the differential equation $(\lambda-A)g=f$ are given by
(3.1) $g(s)=g_{f,K}(s)=\frac{1}{s^{\lambda+1}}\biggl(K-\int_{s}^{1}u^{\lambda}f(u)\,du\biggr),\quad s\in(0,1),\,K\in\mathbb{C}.$
Lemma 3.5.
Let $1\leq p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}.$ Then $\operatorname{Ran}(\lambda-A)=L^{p}[0,1].$
Proof.
On one hand notice that $\frac{1}{s^{\lambda+1}}\in L^{p}[0,1]$ if and only if ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}.$ Therefore, we have to check what happens with the integral part of (3.1). The change of variable $u=e^{t}s$ gives
$\frac{1}{s^{\lambda+1}}\int_{s}^{1}u^{\lambda}f(u)\,du=\int_{0}^{\infty}e^{t(\lambda+1)}(S(t)f)(s)\,dt,\quad s\in(0,1).$
We have already seen in the proof of Theorem 2.2 that $\|S(t)f\|_{p}=e^{-\frac{t}{p}}\|f\|_{p}$. Therefore,
$\int_{0}^{\infty}|e^{t(\lambda+1)}|\|S(t)f\|_{p}\,dt=\|f\|_{p}\int_{0}^{\infty}e^{t({\mathfrak{R}}{\mathfrak{e}}\,\lambda+1-\frac{1}{p})}\,dt,$
where the integral converges for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-\frac{1}{p^{\prime}}$. Thus, $\operatorname{Ran}(\lambda-A)=L^{p}[0,1]$ for all $\lambda\in\mathbb{C}$ such that ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-\frac{1}{p^{\prime}}$ as claimed. ∎
Remark 3.6.
Note that by Proposition 3.2 and Lemma 3.5 the range of $\lambda-A$ is the whole space $L^{p}[0,1]$ ($1\leq p<\infty$) for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\neq-1/p^{\prime}.$ Next result states what happens when ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime}.$
Lemma 3.7.
Let $1\leq p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime}.$ Then $\operatorname{Ran}(\lambda-A)$ is dense on $L^{p}[0,1]$.
Proof.
We will use the fact that the functions with compact support, $C_{c}^{\infty}(0,1)$ are dense in $L^{p}[0,1]$. Let $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime},$ $f\in C_{c}^{\infty}(0,1),$ and define
$g(s)=\frac{1}{s^{\lambda+1}}\biggl(\int_{0}^{1}u^{\lambda}f(u)-\int_{s}^{1}u^{\lambda}f(u)\,du\biggr),\quad s\in(0,1).$
Notice that $g$ is bounded, and therefore $g\in L^{p}[0,1].$ Moreover, by (3.1) $g$ is a solution of the differential equation $(\lambda-A)g=f$ with $g\in D(A).$ This implies that $C_{c}^{\infty}(0,1)\subseteq Ran(\lambda-A)$. As $C_{c}^{\infty}(0,1)$ is dense in $L^{p}[0,1]$ we have that $Ran(\lambda-A)$ is dense as well in $L^{p}[0,1]$. ∎
Now we focus on the study of the range space of the operator $\lambda-B:D(B)\to L^{p}[0,1]$ for a fixed $\lambda\in\mathbb{C}.$ Given $f\in L^{p}[0,1],$ the solutions of the differential equation $(\lambda-B)h=f$ are given by
(3.2) $h(s)=h_{f,K}(s)=s^{\lambda}\biggl(K+\int_{s}^{1}\frac{f(u)}{u^{\lambda+1}}\,du\biggr),\quad s\in(0,1),\,K\in\mathbb{C}.\\$
Lemma 3.8.
Let $1\leq p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$. Then, $\operatorname{Ran}(\lambda-B)=\ker L^{\lambda}$. Additionally, for $p=1$ and ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1$, we have $\operatorname{Ran}(\lambda-B)\subseteq\ker L^{\lambda}$.
Proof.
We first start proving that $\operatorname{Ran}(\lambda-B)\subseteq\ker L^{\lambda}$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$ with $1<p<\infty,$ and also for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1$ if $p=1.$ The idea will be to take a function in $\operatorname{Ran}(\lambda-B)$ that is not in $\ker L^{\lambda}$, and we will reach a contradiction.
First note that the mappings,
$I_{1}:[0,1]$ $\to\mathbb{C}$
$s$ $\mapsto\int_{s}^{1}\frac{f(u)}{u^{\lambda+1}}\,du,\quad f\in L^{p}[0,1],$
and
$I_{2}:[0,1]$ $\to\mathbb{C}$
$s$ $\mapsto\int_{0}^{s}\frac{f(u)}{u^{\lambda+1}}\,du,\quad f\in L^{p}[0,1],$
are continuous by under some assumptions on the value $\lambda,$ because Lemma 2.6 allows to apply the Dominated Convergence Theorem.
Let now $f\in\operatorname{Ran}(\lambda-B)$ with $f\notin\ker L^{\lambda}$ and let $h$ given by (3.2). We have two different scenarios, $K=-L^{\lambda}f$ or $K\neq-L^{\lambda}f$. In the first case, the function $h$ simplifies to,
$h(s)=-s^{\lambda}\int_{0}^{s}\frac{f(u)}{u^{\lambda+1}}\,du,$
and by the continuity of $I_{2}$ one gets $\lim_{s\to 1^{-}}h(s)=-L^{\lambda}f\neq 0$ since $f\notin\ker L^{\lambda},$ hence $h\notin D(B)$, which is a contradiction. On the second scenario, by the continuity of $I_{1}$ we have $\lim_{s\to 0^{+}}h(s)=\lim_{s\to 0^{+}}cs^{\lambda}$ for some complex constant $c\neq 0$. Since $s^{\lambda}\notin L^{p}[0,1]$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$ with $1<p<\infty,$ and also for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1$ if $p=1,$ then $h\notin L^{p}[0,1].$ So, we have reached a contradiction in either case, and we have proved that $\operatorname{Ran}(\lambda-B)\subseteq\ker L^{\lambda}.$
In the following lines we prove that $\ker L^{\lambda}\subseteq\operatorname{Ran}(\lambda-B)$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$ with $1\leq p<\infty$. Let $f\in\ker L^{\lambda}.$ We have that $f$ lies in $\operatorname{Ran}(\lambda-B)$ if and only if there exists $K\in\mathbb{C}$ such that $h$ given by (3.2) is on $L^{p}[0,1]$. Taking $K=L^{\lambda}f=0$ as before, we have
$h(s)=-s^{\lambda}\int_{0}^{s}\frac{f(u)}{u^{\lambda+1}}\,du.$
By a change of variable,
$h(s)=-s^{\lambda}\int_{0}^{s}\frac{f(u)}{u^{\lambda+1}}\,du=-\int_{0}^{\infty}e^{\lambda t}f(e^{-t}s)\,dt=-\int_{0}^{\infty}e^{(\lambda+1)t}(T(t)f)(s)\,dt.$
Thus, applying Minkowski’s inequality and the fact that $\left\|T(t)\right\|=e^{-t/p^{\prime}}$ we get
$\|h\|_{p}\leq\int_{0}^{\infty}e^{({\mathfrak{R}}{\mathfrak{e}}\,\lambda+\frac{1}{p})t}\,dt,$
which is convergent for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-\frac{1}{p}$, so $h\in L^{p}[0,1]$. Then $f\in\operatorname{Ran}(\lambda-B),$ and therefore $\operatorname{Ran}(\lambda-B)=\ker L^{\lambda}$. ∎
Lemma 3.9.
Let $1\leq p<\infty$ and $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p,$ then the $\operatorname{Ran}(\lambda-B)$ has the following density properties
  1. (i)
    if $p=1$, $\operatorname{Ran}(\lambda-B)$ is not dense on $L^{1}[0,1]$,
  2. (ii)
    and if $p>1$, $\operatorname{Ran}(\lambda-B)$ is dense on $L^{p}[0,1]$ when ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p$ and is not dense when ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$.
Proof.
For the case $p=1,$ notice that by Lemma 3.8 $\operatorname{Ran}(\lambda-B)\subseteq\ker L^{\lambda}$, where $\ker L^{\lambda}$ is closed since $L^{\lambda}$ is a bounded operator on $L^{p}[0,1]$ (see Lemma 2.6). Moreover, as $\ker L^{\lambda}\neq L^{1}[0,1]$, see Remark 2.8, we have that
$\overline{\operatorname{Ran}(\lambda-B)}\subseteq\ker L^{\lambda}\neq L^{1}[0,1],$
so $\operatorname{Ran}(\lambda-B)$ is not dense.
On the other hand, when $p>1$ the same idea works when ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$ to prove $\operatorname{Ran}(\lambda-B)$ is not dense on $L^{p}[0,1]$. When ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p,$ the proof of Lemma 2.7 shows that the linear functional $L^{\lambda}$ is defined on $C_{c}^{\infty}(0,1).$ Take $f\in C_{c}^{\infty}(0,1)$ such that $L^{\lambda}f=0.$ We construct the function $h(s)=h_{f,0}(s)$ as in (3.2), that is,
$h(s)=s^{\lambda}\int_{s}^{1}\frac{f(u)}{u^{\lambda+1}}\,du,\quad s\in(0,1).$
Notice that $h\in L^{p}[0,1]$ as it is a continuous function with compact support on $(0,1).$ Then the kernel of the functional $L^{\lambda}$ on $C_{c}^{\infty}(0,1)$ is contained on $\operatorname{Ran}(\lambda-B),$ and also is dense in $L^{p}[0,1]$ (see Corollary 2.9). Therefore, the result follows. ∎
Theorem 3.10.
Let $1\leq p<\infty.$ The differential operator AA satisfies that
  1. (i)
    $\sigma(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}$.
  2. (ii)
    $\sigma_{point}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p^{\prime}\}$.
  3. (iii)
    $\sigma_{ap}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p^{\prime}\}$.
  4. (iv)
    $\sigma_{r}(A)=\emptyset$.
  5. (v)
    $\sigma_{ess}(A)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime}\}$.
Proof.
Items (i) and (ii) are Propositions 3.2 and 3.1, respectively.
Item (iii) follows from the fact that $\sigma_{point}(A)\subseteq\sigma_{ap}(A)$ and $\partial\sigma(A)\subseteq\sigma_{ap}(A),$ and item (iv) follows from Lemma 3.5 and Lemma 3.7.
Lastly, by Proposition 3.1 and Lemma 3.5 we have $\sigma_{ess}(A)\subseteq\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime}\}.$ Moreover, by Proposition 3.2, if $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p^{\prime},$ then it is an accumulation point of both the resolvent set $\rho(A)$ and the spectrum $\sigma(A),$ therefore $\lambda\in\sigma_{ess}(A)$ (see [22, Th. I.3.25]), and the result for the essential spectrum follows. ∎
Theorem 3.11.
Let $1\leq p<\infty.$ The differential operator $B$ satisfies that
  1. (i)
    $\sigma(B)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1/p\}.$
  2. (ii)
    $\sigma_{point}(B)=\emptyset.$
  3. (iii)
    $\sigma_{ap}(B)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p\}.$
  4. (iv)
    For $p=1$, $\sigma_{r}(B)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1\},$
    for $p>1$, $\sigma_{r}(B)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p\}.$
  5. (v)
    $\sigma_{ess}(B)=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p\}$.
Proof.
Item (ii) and (iv) are direct consequences of Proposition 3.1 and Lemma 3.9, and therefore item (i) follows by Proposition 3.3 and the fact that the spectrum is closed.
For item (iii) we have that $\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda=-1/p\}\subseteq\sigma_{ap}(B)$ as the boundary of the spectrum is contained in the approximate point spectrum. Now, if $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$ we have that $(\lambda-B)$ is injective (see Proposition 3.1) and $\operatorname{Ran}(\lambda-B)=\ker L^{\lambda}$ (see Lemma 3.8) which is closed, therefore $\lambda\notin\sigma_{ap}(B)$ and the result follows.
Finally, for item (v), let $\lambda\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/p$. By the Banach isomorphism theorem [33, Theorem 1.7.14] and Lemma 3.8 we have that
${\raisebox{1.99997pt}{$L^{p}[0,1]$}\left/\raisebox{-1.99997pt}{$\operatorname{Ran}(\lambda-B)$}\right.}={\raisebox{1.99997pt}{$L^{p}[0,1]$}\left/\raisebox{-1.99997pt}{$\ker L^{\lambda}$}\right.}\cong\operatorname{Ran}L^{\lambda}.$
Therefore, $\operatorname{codim}(\lambda-B)=\dim\operatorname{Ran}L^{\lambda}=1$, because $L^{\lambda}$ is the non-zero linear functional. Moreover, by Proposition 3.1 we have that $\dim\ker(\lambda-B)=0,$ therefore $\lambda\notin\sigma_{ess}(B)$. Finally, by [22, Th. I.3.25] the elements which are accumulation points of the spectrum and the resolvent set at the same time belong to the essential spectrum, so item (v) follows. ∎
4. Spectra of Koopman semigroups
At Section 2 we introduced the operator families $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ which are $C_{0}$-semigroups for $1\leq p<\infty,$ and which have the operators AA and $B$ as infinitesimal generators respectively. In this section, we use the spectral study of the generators in Section 3 in order to study the spectral properties of the $C_{0}$-semigroups $(T(t))_{t\geq 0}$ and $(S(t))_{t\geq 0}$ defined on $L^{p}[0,1]$ for $1\leq p<\infty$. From now on, we denote by $B(c,r)$ the open ball with center $c$ and radius $r$. Also, from now on, the following subspaces
$M_{p,r}:=\{f\in L^{p}[0,1]\,:\,f(s)=0\text{ a.e. for }s\in(0,r)\},\quad r\in(0,1),\,1\leq p\leq\infty,$
will be play a key role in some main results in this paper.
Theorem 4.1.
Let $1\leq p\leq\infty$ and $t>0$. The bounded operator $T(t)$ satisfies that
  1. (i)
    $\sigma(T(t))=\overline{B(0,e^{-t/p^{\prime}})}$.
  2. (ii)
    For $1\leq p<\infty$, $\sigma_{point}(T(t))=B(0,e^{-t/p^{\prime}})$,
    for $p=\infty$, $\sigma_{point}(T(t))=\overline{B(0,e^{-t})}.$
  3. (iii)
    $\sigma_{ap}(T(t))=\overline{B(0,e^{-t/p^{\prime}})}$.
  4. (iv)
    $\sigma_{r}(T(t))=\emptyset$ for $1\leq p<\infty.$
  5. (v)
    $\sigma_{ess}(T(t))=\overline{B(0,e^{-t/p^{\prime}})}$.
Proof.
First, we prove the results for the case when $1\leq p<\infty$, that is, when $(T(t))_{t\geq 0}$ is a strongly continuous semigroup.
For item (i), we have by [23, Theorem IV.3.6] that
$e^{t\sigma(A)}=\overline{B(0,e^{-t/p^{\prime}})}\setminus\{0\}\subseteq\sigma(T(t))$
and as the growth bound of $T(t)$ is $w_{0}=-1/p^{\prime}$, we have that $r(T(t))=e^{-t/p^{\prime}}.$ Then $\sigma(T(t))\subseteq\overline{B(0,e^{-t/p^{\prime}})}$ and taking closures we get the desired result.
From [23, Theorem IV.3.7], we have that
$\sigma_{point}(T(t))\setminus\{0\}=B(0,e^{-t/p^{\prime}})\setminus\{0\}.$
Moreover, $0\in\sigma_{point}(T(t))$ and item (ii) follows. Indeed, take $f=\chi_{(e^{-t},1]}$, then $T(t)f(s)=0$ with $s\in[0,1].$
Analogously, we have by [23, Theorem IV.3.7] that,
$\sigma_{r}(T(t))\setminus\{0\}=\emptyset.$
Even more, let $f\in L^{p}[0,1]$ and define $g(s)=e^{t}\chi_{[0,e^{-t})}(s)f(e^{t}s)$. It’s easy to check that under a change of variables $\left\|g\right\|_{p}=\left\|f\right\|_{p}<\infty$, and $T(t)g(s)=f(s)$, so $\operatorname{Ran}(T(t))=L^{p}[0,1]$. Therefore, the range is dense and $0\notin\sigma_{r}(T(t))$ and item (iv) follows.
For the approximate point spectrum we will use the fact that $\sigma(T(t))=\sigma_{ap}(T(t))\cup\sigma_{r}(T(t))$, then the result follows as $\sigma_{r}(T(t))=\emptyset$.
Finally, for item (v), if $\mu\in\sigma(T(t))\setminus\{0\}$ with $\mu=e^{\lambda t}$ with $\lambda\in\sigma_{point}(A)$ we have that the function $\frac{1}{s^{\lambda+1}}$ is an eigenfunction for $\mu$. Thus, every eigenvalue $\mu$ has infinite eigenfunctions associated with it (with periodicity $2\pi$) and $\ker(\mu-T(t))$ is infinite dimensional so $\mu\in\sigma_{ess}(T(t))$. On the other hand, notice that $\ker(T(t))=M_{p,{e^{-t}}}.$ Then $\dim(\ker(T(t)))=\infty$ and $0\in\sigma_{ess}(T(t))$. Lastly, by [22, Th. I.3.25] the elements which are accumulation points of $\sigma(T(t))$ and $\rho(T(t))$ at the same time belong to $\sigma_{ess}(T(t))$ and result follows.
For $p=\infty$ we use the fact that the operator $T(t)$ defined on $L^{\infty}[0,1]$ is the dual operator of $S(t)$ defined on $L^{1}[0,1]$, therefore the results for items (i), (ii) follow from Theorem 4.2 below and [23, Corollary B.12, Proposition IV.1.12]. Item (iii) follows since $\sigma_{point}(T(t))\subseteq\sigma_{ap}(T(t))\subseteq\sigma(T(t)).$ For item (v) we divide our study into 3 different cases. First the values $\lambda\in\mathbb{C}$ which are in the boundary of the spectrum, that is, $|\lambda|=e^{-t},$ because they are accumulation points of both the spectrum and resolvent set, and therefore they belong to $\sigma_{ess}(T(t))$ (see [22, Th. I.3.25]). Additionally, $0\in\sigma_{ess}(T(t))$ because $\ker(T(t))=M_{\infty,e^{-t}}$ which is infinite dimensional. Lastly, take $\lambda\in B(0,e^{-t})\setminus\{0\},$ then $\lambda=e^{\mu t}$ with $\mu\in\mathbb{C}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\mu<-1$. Although in this case $T(t)$ doesn’t define a strongly continuous semigroup, we still have that for $n\in\mathbb{Z}$ the functions $g_{\lambda,n}(s):=\frac{1}{s^{\mu+\frac{2\pi in}{t}+1}}\in L^{\infty}[0,1]$ are eigenfunctions for the eigenvalue $\lambda$. Therefore, $\lambda\in\sigma_{ess}(T(t))$ since $\dim\ker(\lambda-T(t))=\infty$. ∎
Theorem 4.2.
Let $1\leq p\leq\infty$ and $t>0$. The bounded operator $S(t)$ satisfies that
  1. (i)
    $\sigma(S(t))=\overline{B(0,e^{-t/p})}$.
  2. (ii)
    $\sigma_{point}(S(t))=\emptyset$.
  3. (iii)
    $\sigma_{ap}(S(t))=\delta B(0,e^{-t/p})$.
  4. (iv)
    For $p=1$, $\sigma_{r}(S(t))=\overline{B(0,e^{-t})},$
    for $1<p<\infty$, $\sigma_{r}(S(t))=B(0,e^{-t/p}),$
    and for $p=\infty$, $B(0,e^{-t/p})\subseteq\sigma_{r}(S(t)).$
  5. (v)
    $\sigma_{ess}(S(t))=\overline{B(0,e^{-t/p})}$.
Proof.
As in the last theorem, we prove the results for the case when $1\leq p<\infty$, that is, when $(S(t))_{t\geq 0}$ is a strongly continuous semigroup.
Item (i) follows the same idea as in Theorem 4.1, and for item (ii) we use [23, Theorem IV.3.7] for the point spectrum and a direct computation shows that $S(t)$ is injective.
For the residual spectrum, we have that, if $p=1$
$\sigma_{r}(S(t))\setminus\{0\}=\overline{B(0,e^{-t/p})}\setminus\{0\},$
and for $p>1$,
$\sigma_{r}(S(t))\setminus\{0\}=B(0,e^{-t/p})\setminus\{0\}.$
Moreover, notice that $0\in\sigma_{r}(S(t))$ in both cases ($p=1$ and $p>1$) since $\operatorname{Ran}(S(t))=M_{p,e^{-t}}^{\perp}=\{f\in L^{p}(0,1):f(x)=0\text{ a.e. for }x\in(e^{-t},1)\},$ which is a closed subspace of $L^{p}[0,1]$ and therefore not dense.
Next, we prove item (iii) using the characterization of the approximate point spectrum via sequences (see [23, Lemma IV.1.9]). Take $\lambda\in B(0,e^{-t/p})\setminus\{0\}$ and suppose $\lambda\in\sigma_{ap}(S(t))$, then there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subset L^{p}[0,1]$, such that $\left\|f_{n}\right\|_{p}=1$ and $\lim_{n\to\infty}\left\|S(t)f_{n}-\lambda f_{n}\right\|_{p}=0$. Therefore,
$\left\|S(t)f_{n}-\lambda f_{n}\right\|_{p}^{p}=\int_{0}^{e^{-t}}\left|f_{n}(e^{t}u)-\lambda f_{n}(u)\right|^{p}\,du+\int_{e^{-t}}^{1}\left|\lambda f_{n}(u)\right|^{p}\,du\to 0$
as $n\to\infty$. As both integrals are positive this means that each of them go to zero. However, doing a change of variables the first integral becomes
$\int_{0}^{e^{-t}}\left|f_{n}(e^{t}u)-\lambda f_{n}(u)\right|^{p}\,du$ $=\int_{0}^{1}\left|f_{n}(x)-\lambda f_{n}(e^{-t}x)\right|^{p}e^{-t}\,dx$
$=\left|\lambda\right|^{p}e^{(p-1)t}\int_{0}^{1}\left|T(t)f_{n}(x)-\frac{e^{-t}}{\lambda}f_{n}(x)\right|^{p}\,dx\to 0.$
This implies that the value $\frac{e^{-t}}{\lambda}\in\sigma_{ap}(T(t))$ on the space $L^{p}[0,1]$. But, notice that $\left|\frac{e^{-t}}{\lambda}\right|>e^{-t}e^{t/p}=e^{-t/p^{\prime}}$ as $\lambda\in B(0,e^{-t/p})\setminus\{0\}$, which is a contradiction as $\sigma_{ap}(T(t))=\overline{B(0,e^{-t/p^{\prime}})}$. The fact that $0\notin\sigma_{ap}(S(t))$ follows from the fact that $\operatorname{Ran}(S(t))=\{f\in L^{p}[0,1]\,:\,f(s)=0\text{ a.e. for }s\in(e^{-t},1)\}$ is closed. This case finishes with the fact that the boundary of the spectrum is always included in the approximate point spectrum.
For item (v) we use the duality of the operators $T(t)$ and $S(t)$ and the properties of the essential spectrum and duality, see [22, Th. IX.1.1].
The case $p=\infty$. The operator $S(t)$ defined on $L^{\infty}[0,1]$ is the dual of $T(t)$ defined on $L^{1}[0,1]$, therefore the results for items (i) and (ii) follow from Theorem 4.1 and [23, Corollary B.12, Proposition IV.1.12]. Also, by the last paragraph we have also the result for item (v). The proof for the approximate spectrum is similar that for the case $1\leq p<\infty$ using the $\|\cdot\|_{\infty}$ norm. Finally, item (iv) follows by $\sigma(S(t))=\sigma_{ap}(S(t))\cup\sigma_{r}(S(t)).$
5. The fine spectra of the generalized Cesàro-Hardy operators
In this section, we apply the results obtained in the preceding sections to study the boundedness and the spectrum on $L^{p}[0,1]$ of the Cesàro-Hardy operator $\mathcal{C}_{\alpha}$ and its dual $\mathcal{C}_{\alpha}^{*}$. In order to do so, we will use the theory of the natural functional calculus for sectorial operators defined by Haase in [27], but first we must fix some notation.
The open sector of angle $\omega$ is defined as $S_{\omega}:=\{z\in\mathbb{C}\,:\,z\neq 0\text{ and }\left|\arg z\right|<\omega\}$, for $0<\omega\leq\pi$ and $S_{0}:=(0,\infty)$. Let $0\leq\omega<\pi$, an operator $\mathcal{A}$ on a Banach space $X$ is called sectorial of angle $\omega$ if $\sigma(\mathcal{A})\subset\overline{S_{\omega}}$ and $\sup\{\left\|\lambda(\lambda-\mathcal{A})^{-1}\right\|\,:\,\lambda\notin\overline{S_{\omega^{\prime}}}\}<\infty$ for all $\omega<\omega^{\prime}<\pi$. In particular if $\mathcal{A}$ generates a uniformly bounded semigroup $(\mathcal{T}(t))_{t\geq 0}$, i.e. $\left\|\mathcal{T}(t)\right\|\leq C$ for some positive constant $C$, $\mathcal{A}$ is sectorial of angle $\pi/2$ by the Hille-Yosida Theorem ([23, Corollary II.3.6]). Now we are prepared to prove the results for Cesàro-Hardy operators.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$, we consider the Cesàro-Hardy operator of order $\alpha$ given by,
$\mathcal{C}_{\alpha}f(s)=\frac{\alpha}{s^{\alpha}}\int_{0}^{s}(s-u)^{\alpha-1}f(u)\,du,\quad s\in(0,1),$
and the dual Cesàro-Hardy operator of order $\alpha$ given by
$\mathcal{C}_{\alpha}^{*}f(s)=\alpha\int_{s}^{1}\frac{(u-s)^{\alpha-1}}{u^{\alpha}}f(u)\,du,\quad s\in(0,1).$
Theorem 5.1.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$. Then:
  1. (i)
    The operator $\mathcal{C}_{\alpha}$ is a bounded operator on $L^{p}[0,1]$, for $1<p\leq\infty$.
  2. (ii)
    The operator $\mathcal{C}_{\alpha}^{*}$ is a bounded operator on $L^{p}[0,1]$, for $1\leq p<\infty$.
In addition, for $f\in L^{p}[0,1]$ the following subordination identities hold,
$\mathcal{C}_{\alpha}f(s)=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}T(t)f(s)\,dt,\quad s\in[0,1],\,1<p\leq\infty,$
and
$\mathcal{C}_{\alpha}^{*}f(s)=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}S(t)f(s)\,dt,\quad s\in[0,1],\,1\leq p<\infty.$
Proof.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$ and $f\in L^{p}[0,1]$. Then, the change of variable $u=e^{-t}s$ yields,
$\mathcal{C}_{\alpha}f(s)=\frac{\alpha}{s^{\alpha}}\int_{0}^{s}(s-u)^{\alpha-1}f(u)\,du=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}T(t)f(s)\,dt,$
and, the change of variable $u=e^{t}s$ gives,
$\mathcal{C}_{\alpha}^{*}f(s)=\alpha\int_{s}^{1}\frac{(u-s)^{\alpha-1}}{u^{\alpha}}f(u)\,du=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}S(t)f(s)\,dt.$
Using the boundedness of the operators $T(t)$ and $S(t)$ we get
$\left\|\mathcal{C}_{\alpha}f\right\|_{p}\leq\left|\alpha\right|\left\|f\right\|_{p}\int_{0}^{\infty}(1-e^{-t})^{{\mathfrak{R}}{\mathfrak{e}}\,\alpha-1}e^{-t/p^{\prime}}\,dt=\left|\alpha\right|\left\|f\right\|_{p}\frac{\Gamma({\mathfrak{R}}{\mathfrak{e}}\,\alpha)\Gamma(1/p^{\prime})}{\Gamma({\mathfrak{R}}{\mathfrak{e}}\,\alpha+1/p^{\prime})},\quad 1<p\leq\infty,$
and
$\left\|\mathcal{C}_{\alpha}^{*}f\right\|_{p}\leq\left|\alpha\right|\left\|f\right\|_{p}\int_{0}^{\infty}(1-e^{-t})^{{\mathfrak{R}}{\mathfrak{e}}\,\alpha-1}e^{-t/p}\,dt=\left|\alpha\right|\left\|f\right\|_{p}\frac{\Gamma({\mathfrak{R}}{\mathfrak{e}}\,\alpha)\Gamma(1/p)}{\Gamma({\mathfrak{R}}{\mathfrak{e}}\,\alpha+1/p)},\quad 1\leq p<\infty.\qed$
The following remark outlines the duality between the generalized Cesàro-Hardy operators. The proof is straightforward and is therefore omitted.
Remark 5.2.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$. The Cesàro-Hardy operators $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\alpha}^{*}$ show a dual relationship as follows:
  • For $1\leq p<\infty$, $\mathcal{C}_{\alpha}^{*}$ on $L^{p^{\prime}}[0,1]$ is the dual operator of $\mathcal{C}_{\alpha}$ on $L^{p}[0,1]$.
  • For $1<p\leq\infty$, $\mathcal{C}_{\alpha}$ on $L^{p}[0,1]$ is the dual operator of $\mathcal{C}_{\alpha}^{*}$ on $L^{p^{\prime}}[0,1]$.
Theorem 5.3.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$ and $1\leq p<\infty$. Then the operator $\mathcal{C}_{\alpha}^{*}:L^{p}[0,1]\rightarrow L^{p}[0,1]$ satisfies
  1. (i)
    $\sigma(\mathcal{C}_{\alpha}^{*})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\}.$
  2. (ii)
    $\sigma_{point}(\mathcal{C}_{\alpha}^{*})=\emptyset.$
  3. (iii)
    $\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in i\mathbb{R}\right\}\cup\{0\}\subset\sigma_{ap}(\mathcal{C}_{\alpha}^{*})$.
  4. (iv)
    For $p=1$, $\sigma_{r}(\mathcal{C}_{\alpha}^{*})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+1)}{\Gamma(\alpha+z+1)}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}$,
    a
    and for $1<p<\infty$, $\sigma_{r}(\mathcal{C}_{\alpha}^{*})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in\mathbb{C}_{+}\right\}$.
  5. (v)
    $\sigma_{ess}(\mathcal{C}_{\alpha}^{*})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in i\mathbb{R}\right\}\cup\{0\}.$
Proof.
For $1\leq p<\infty$, the family $(e^{t/p}S(t))_{t\geq 0}$ forms a uniformly bounded $C_{0}$-semigroup on $L^{p}[0,1]$ with generator $(B+1/p,D(B))$ (see Theorem 2.2). On the other hand, let $\tilde{\mu}$ be the finite Borel measure on $0,\infty)$ and absolutely continuous with respect to the Lebesgue measure, defined by $d\tilde{\mu}(t)=\alpha e^{-t/p}(1-e^{-t})^{\alpha-1}dt$. Notice that the Laplace transform of the measure $\tilde{\mu}$ is,
$\mathcal{L}(\tilde{\mu})(z):=\int_{[0,\infty)}e^{-zt}\mu(dt)=\int_{0}^{\infty}\alpha e^{-t(z+1/p)}(1-e^{-t})^{\alpha-1}dt=\frac{\Gamma(\alpha+1)\Gamma(z+1/p)}{\Gamma(\alpha+z+1/p)},$
[…]
(5.1) $\frac{\Gamma(z+\alpha)}{\Gamma(z)}=z^{\alpha}(1+\frac{\alpha(\alpha+1)}{2z}+O(\left|z\right|^{-2})),\quad\left|z\right|\to\infty,$
[…]
$\sigma(\mathcal{C}_{\alpha}^{*})=\mathcal{L}(\tilde{\mu})(\sigma(-B-1/p))\cup\{0\}=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\}.$
Item (iv) follows from the fact that the dual operator of $\mathcal{C}_{\alpha}^{*}$ defined on $L^{p}[0,1]$ with $1\leq p<\infty$ is $\mathcal{C}_{\alpha}$ defined on $L^{p^{\prime}}[0,1]$. For a closed densely defined operator, its residual operator coincides with the point spectrum of its dual, see [23, Proposition IV.1.12] and thus item (ii) of Theorem 5.4 below finishes the proof of this item.
Item (v) follows from the spectral mapping theorem for the essential spectrum that appears in [37, Theorem 5.4],
$\sigma_{ess}(\mathcal{C}_{\alpha}^{*})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in i\mathbb{R}\right\}\cup\{0\},$
and by (5.1) and [37, Theorem 5.6] we have that,
$\sigma_{point}(\mathcal{C}_{\alpha}^{*})=\mathcal{L}(\tilde{\mu})(\sigma_{point}(-B-1/p))=\emptyset.$
In [37, Theorem 5.4 and Theorem 5.6] the author uses the notion of quasi-regularity; notice that the property of having polynomial limit is stronger, so we are under the assumptions of those results.
Lastly, item (iii) follows from the fact that the boundary of the spectrum is always contained in the approximate spectrum and the fact that $\sigma_{ap}(\mathcal{C}_{\alpha}^{*})\cup\sigma_{r}(\mathcal{C}_{\alpha}^{*})=\sigma(\mathcal{C}_{\alpha}^{*})$. ∎
Theorem 5.4.
Let ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$ and $1<p\leq\infty$. Then the operator $\mathcal{C}_{\alpha}:L^{p}[0,1]\rightarrow L^{p}[0,1]$ satisfies
  1. (i)
    $\sigma(\mathcal{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p^{\prime}})}{\Gamma(\alpha+z+\frac{1}{p^{\prime}})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\}.$
  2. (ii)
    For $1<p<\infty$, $\sigma_{point}(\mathcal{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p^{\prime}})}{\Gamma(\alpha+z+\frac{1}{p^{\prime}})}\,:\,z\in\mathbb{C}_{+}\right\},$
    a
    and for $p=\infty$, $\sigma_{point}(\mathcal{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+1)}{\Gamma(\alpha+z+1)}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}$.
  3. (iii)
    $\sigma_{ap}(\mathcal{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p^{\prime}})}{\Gamma(\alpha+z+\frac{1}{p^{\prime}})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\}.$
  4. (iv)
    For $1<p<\infty$, $\sigma_{r}(\mathcal{C}_{\alpha})=\emptyset$.
  5. (v)
    $\sigma_{ess}(\mathcal{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p^{\prime}})}{\Gamma(\alpha+z+\frac{1}{p^{\prime}})}\,:\,z\in i\mathbb{R}\right\}\cup\{0\}$.
Proof.
Item (i) and (v) follow from the duality property stated in Remark 5.2 and the duality of the spectrum and essential spectrum, see [22, Th. IX.1.1].
When $1<p<\infty$ the result for the point spectrum follows the same ideas as the proof of the point spectrum for the dual Cesàro-Hardy operator. For the $p=\infty$ case notice that by the result for the point spectrum for the family $(T(t))_{t\geq 0},$ we have that $g_{\lambda}(s)=\frac{1}{s^{\lambda+1}}$ is an eigenfunction with eigenvalue $e^{t\lambda}$ with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1$, therefore,
$\mathcal{C}_{\alpha}g_{\lambda}=\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}T(t)g_{\lambda}\,dt=g_{\lambda}\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}e^{t\lambda}\,dt=\frac{\Gamma(\alpha+1)\Gamma(-\lambda)}{\Gamma(\alpha-\lambda)}g_{\lambda},$
with ${\mathfrak{R}}{\mathfrak{e}}\,\lambda\leq-1$, therefore
$\left\{\frac{\Gamma(\alpha+1)\Gamma(z+1)}{\Gamma(\alpha+z+1)}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\subseteq\sigma_{point}(\mathcal{C}_{\alpha}).$
Moreover, $0\notin\sigma_{\text{point}}(\mathcal{C}_{\alpha})$ by a theorem of E.C. Titchmarsh [43, Theorem VII], which states that if $f,g\in L^{1}[0,1]$ satisfy
$\int_{0}^{s}f(u-s)g(u)\,du=0\quad\text{for almost every }s\in(0,1),$
and $f\neq 0$ almost everywhere on $[0,1]$, then $g$ must be zero almost everywhere on $[0,1]$. As $\mathcal{C}_{\alpha}(s)=\frac{\alpha}{s^{\alpha}}\int_{0}^{s}(s-u)^{\alpha-1}f(u)\,du$, the last theorem implies that if we have $\mathcal{C}_{\alpha}f=0$ then $f=0$ so $\mathcal{C}_{\alpha}$ is injective. As $\sigma_{point}(\mathcal{C}_{\alpha})\subseteq\sigma(\mathcal{C}_{\alpha})$ the result follows.
Item (iv) follows the same argument as in the proof of Theorem 5.3, although in this case the proof doesn’t apply for the case $p=\infty$.
Lastly, for item (iii), notice that when $1<p<\infty$ by $\sigma_{ap}(\mathcal{C}_{\alpha})\cup\sigma_{r}(\mathcal{C}_{\alpha})=\sigma(\mathcal{C}_{\alpha}),$ item (i) and (iv), the result follows. For $p=\infty$, notice that $0\in\sigma_{ap}(\mathcal{C}_{\alpha})$. Indeed, consider the sequence $f_{n}(s)=s^{n}$, $s\in(0,1)$ and $n\in\mathbb{N}$. Thus,
$\left|\mathcal{C}_{\alpha}f_{n}(s)\right|\leq\left|\alpha\right|s^{n}\int_{0}^{\infty}(1-e^{-t})^{{\mathfrak{R}}{\mathfrak{e}}\,\alpha-1}e^{-(n+1)t}\,dt\to 0,\quad\text{as }n\to\infty$
by the dominated convergence theorem. Using the characterization of the approximate point spectrum via sequences (see [23, Lemma IV.1.9]) we have that $0\in\sigma_{ap}(\mathcal{C}_{\alpha})$. Also, by definition, $\sigma_{point}(\mathcal{C}_{\alpha})\subseteq\sigma_{ap}(\mathcal{C}_{\alpha}),$ so the result follows. ∎
Remark $5.$
Note that in particular the classical Cesàro-Hardy operator $\mathcal{C}:=\mathcal{C}_{1}$ is the inverse of the differential operator A-A (see Equation 2.1 and Proposition 2.5), that is,
$\mathcal{C}f(s)=\frac{1}{s}\int_{0}^{s}f(u)\,du=\int_{0}^{\infty}(T(t)f)(s)\,dt=(-A)^{-1}f(s),\quad s\in(0,1),\,f\in L^{p}[0,1].$
6. Connection with discrete generalized Cesàro operators on $\ell^{p}$
In this section we consider the usual sequence Lebesgue space $\ell^{p}$ with the norm
$\left(\sum_{n=0}^{\infty}|a(n)|^{p}\right)^{\frac{1}{p}},$
for $1\leq p<\infty,$ and $\sup_{n}|a(n)|<\infty$ for $p=\infty.$ Let us denote by $\mathscr{C}_{\alpha}$ and $\mathscr{C}_{\alpha}^{*}$ the generalized discrete Cesàro operator for ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$ and its dual operator, given by
$\mathscr{C}_{\alpha}a(n)$ $:=$ $\frac{1}{k^{\alpha+1}(n)}\sum_{j=0}^{n}k^{\alpha}(n-j)a(j),$
$\mathscr{C}^{*}_{\alpha}a(n)$ $:=$ $\sum_{j=n}^{\infty}\frac{1}{k^{\alpha+1}(j)}k^{\alpha}(j-n)a(j),$
for $a\in\ell^{p},$ and $k^{\alpha}(n):=\displaystyle{\frac{\Gamma(n+\alpha)}{\Gamma(\alpha)\Gamma(n+1)}}$ which are the classical Cesàro numbers for $n\in\mathbb{N}_{0}$. Note that $\mathscr{C}_{1}=C_{0}.$
The aim of this section is to present a nice connection between $\mathcal{C_{\alpha}}$ and $\mathscr{C}_{\alpha}$ and then $\mathcal{C}_{\alpha}^{*}$ and $\mathscr{C}_{\alpha}^{*}$. In the next theorem, we present a way to do this. The proof of the following theorem straightforward and is left to the reader.
Theorem 6.1.
Take $1\leq p\leq\infty$.
  • (i)
    The operator $V:L^{p}[0,1]\to\ell^{p}$ given by
    $V(f)(n):=\int_{0}^{1}{\frac{t^{n}}{n!}}e^{-t}f(t)dt,\qquad n\in\mathbb{N},\quad f\in L^{p}[0,1],$
    is well-defined, linear and $\|V\|\leq 1$.
  • (ii)
    The operator $V^{*}:\ell^{p}\to L^{p}[0,1]$ given by
    $V^{*}(a)(t):=\sum_{n=0}^{\infty}{\frac{t^{n}}{n!}}e^{-t}a(n),\qquad t\in[0,1],\quad a\in\ell^{p},$
    is also well-defined, linear and $\|V\|\leq 1$. In particular, $V^{*}$ on $\ell^{p}$ is the dual operator of $V$ on $L^{p^{\prime}}[0,1]$ for $1<p\leq\infty.$
Remark 6.2.
Note that both operators are not invertible for $1\leq p\leq\infty$. In the case that $V$ was invertible, there will exist $f\in L^{p}[0,1]$ such that $V(f)=\delta_{0}$, where $\delta_{0}(n):=1$ if $n=0$ and $\delta_{0}(n):=0$ for $n>0$. Then
$\sum_{n=0}^{\infty}Vf(n)z^{n}=1=\int_{0}^{1}f(t)e^{-t(1-z)}dt,\qquad z\in\mathbb{C},\,|z|<1.$
This implies that the function $f$ should be the delta Dirac distribution, which is a contradiction.
Now we suppose that $V^{*}$ is invertible for $1\leq p\leq\infty$. Then there exists $a\in\ell^{p}$ such that $V^{*}(a)=\chi_{[0,1]}$, i.e.,
$e^{t}=\sum_{n=0}^{\infty}a(n){\frac{t^{n}}{n!}},\qquad t\in[0,1].$
Then $a(n)=1$ for $n\geq 0$ and we obtain a contradiction $(a\in\ell^{p})$ for $1\leq p<\infty$. For $p=\infty$, and $V^{*}$ is invertible, there exists $b\in\ell^{\infty}$ such that $V^{*}(b)(t)=t$, i.e.,
$te^{t}=\sum_{n=0}^{\infty}b(n){\frac{t^{n}}{n!}},\qquad t\in[0,1],$
and we obtain a contradiction with $b\in\ell^{\infty}$.
Now we consider the one-parameter families of operators $(\mathscr{T}(t))_{t\geq 0}$ and $(\mathscr{S}(t))_{t\geq 0}$ acting on $\ell^{p}$ with $1\leq p\leq\infty$, where
(6.1) $\mathscr{T}(t)a(n)$ $:=$ $\sum_{j=0}^{n}\binom{n}{j}e^{-tj}(1-e^{-t})^{n-j}a(j),$
(6.2) $\mathscr{S}(t)a(n)$ $:=$ $e^{-tn}\sum_{j=n}^{\infty}\binom{j}{n}(1-e^{-t})^{j-n}a(j),$
for $n\in\mathbb{N}_{0}$ and $a\in\ell^{p}$. In fact, both families $(\mathscr{T}(t))_{t\geq 0}$ and $(\mathscr{S}(t))_{t\geq 0}$ are dual $C_{0}$-semigroups in $\ell^{p}$, $\|\mathscr{T}(t)\|\leq e^{\frac{t}{p}}$ and $\|\mathscr{S}(t)\|\leq e^{t(1-{\frac{1}{p}})}$ for $1\leq p\leq\infty$. As we have explained in the introduction, these semigroups were introduced in [1] to study spectral properties of $C_{0}$ and its generalizations $\mathscr{C}_{\alpha}$ and $\mathscr{C}_{\alpha}^{*}$ on $\ell^{p}$ (and also on other sequence spaces).
In the following theorem, we present that the one-parameter families $(T(t))_{t\geq 0}$, $(S(t))_{t\geq 0}$, $(\mathscr{T}(t))_{t\geq 0}$ and $(\mathscr{S}(t))_{t\geq 0}$ intertwines with $V$ and $V^{*}$.
Theorem 6.3.
Let $1\leq p\leq\infty,$ then
  • (i)
    $\mathscr{S}(t)\circ V=e^{t}(V\circ S(t))$ for $t>0$.
  • (ii)
    $V^{*}\circ\mathscr{T}(t)=e^{t}(T(t)\circ V^{*})$ for $t>0$.
Proof.
(i) Take $t>0$ and $f\in L^{p}[0,1]$ with $1\leq p\leq\infty$. For $n\in\mathbb{N}_{0}$, one obtains that
$(\mathscr{S}(t)\circ V)(f)(n)$ $=$ $e^{-tn}\displaystyle\sum_{j=n}^{\infty}\binom{j}{n}(1-e^{-t})^{j-n}\int_{0}^{1}\frac{s^{j}}{j!}e^{-s}f(s)\,ds$
$=$ $e^{-tn}\int_{0}^{1}e^{-s}f(s)\displaystyle\sum_{j=n}^{\infty}\frac{1}{n!(j-n)!}(1-e^{-t})^{j-n}s^{j}\,ds$
$=$ $e^{-tn}\int_{0}^{1}e^{-s}f(s)\frac{s^{n}}{n!}\displaystyle\sum_{m=0}^{\infty}\frac{(s(1-e^{-t}))^{m}}{m!}\,ds$
$=$ $e^{-tn}\int_{0}^{1}e^{-s}f(s)\frac{s^{n}}{n!}e^{s(1-e^{-t})}\,ds$
$=$ $e^{t}\int_{0}^{e^{-t}}f(e^{t}u)\frac{u^{n}}{n!}e^{-u}\,du=e^{t}(V\circ S(t))(f)(n).$
(ii) Now take $t>0$ and $a\in\ell^{p}$ with $1\leq p\leq\infty$. For $0<s<1$, we get
$(V^{*}\circ\mathscr{T}(t))(a)(s)$ $=$ $\sum_{n=0}^{\infty}e^{-s}\frac{s^{n}}{n!}\displaystyle\sum_{j=0}^{n}\binom{n}{j}e^{-tj}(1-e^{-t})^{n-j}a(j)$
$=$ $e^{-s}\sum_{j=0}^{\infty}\frac{e^{-tj}}{j!}a(j)\displaystyle\sum_{n=j}^{\infty}\frac{s^{n}(1-e^{-t})^{n-j}}{(n-j)!}$
$=$ $e^{-s}\sum_{j=0}^{\infty}\frac{(se^{-t})^{j}}{j!}a(j)\displaystyle\sum_{m=0}^{\infty}\frac{(s(1-e^{-t}))^{m}}{m!}\,ds$
$=$ $\sum_{j=0}^{\infty}\frac{(se^{-t})^{j}}{j!}e^{-se^{-t}}a(j)=e^{t}(T(t)\circ V^{*})(a)(s),$
and we conclude the proof. ∎
Remark 6.4.
In the case that $1\leq p<\infty$, we may prove Theorem 6.3 (ii) by duality. For $t>0$ we get
$V^{*}\circ\mathscr{T}(t)=(\mathscr{S}(t)\circ V)^{*}=e^{t}(V\circ S(t))^{*}=e^{t}(T(t)\circ V^{*}).$
The following theorem presents the connection between ${\mathcal{C}}_{\alpha}$, ${\mathcal{C}}_{\alpha}^{*}$, $\mathscr{C}_{\alpha}$ and $\mathscr{C}_{\alpha}^{*}$. Remind that for $a\in\ell^{p}$ and $n\in\mathbb{N}_{0}$, we have that
${\mathscr{C}}_{\alpha}a(n)$ $=$ $\displaystyle\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}e^{-t}\mathscr{T}(t)a(n)\,dt,\quad\,1<p\leq\infty,$
${\mathscr{C}}^{*}_{\alpha}a(n)$ $=$ $\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}e^{-t}\mathscr{S}(t)a(n)\,dt,\quad\,1\leq p<\infty,$
see [1, Theorem 7.1].
Theorem 6.5.
Take ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0.$ Then
  • (i)
    $V\circ{\mathcal{C}}_{\alpha}^{*}=\mathscr{C}^{*}_{\alpha}\circ V$ for $1\leq p<\infty$.
  • (ii)
    $V^{*}\circ\mathscr{C}_{\alpha}={\mathcal{C}}_{\alpha}\circ V$ for $1<p\leq\infty$.
Proof.
(i) Take $f\in L^{p}[0,1]$ for $1\leq p<\infty$. By Theorem 5.1 and Theorem 6.3(i), we obtain that
$(V\circ{\mathcal{C}}_{\alpha}^{*}f)(n)$ $=$ $\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}(V\circ S(t)f)(n)ds$
$=$ $\alpha\int_{0}^{\infty}(1-e^{-t})^{\alpha-1}e^{-t}(\mathscr{S}(t)\circ Vf)(n)ds=(\mathscr{C}^{*}_{\alpha}\circ Vf)(n),$
for $n\in\mathbb{N}_{0}$. Similarly, it is proven the item (ii). ∎
Remark $6.$
Take ${\mathfrak{R}}{\mathfrak{e}}\,\alpha>0$, then
$\sigma(\mathcal{C_{\alpha}})=\sigma(\mathscr{C}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p^{\prime}})}{\Gamma(\alpha+z+\frac{1}{p^{\prime}})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\},$
for $1<p\leq\infty,$ see Theorem 5.4 and [1, Theorem 7.6], and
$\sigma(\mathcal{C^{*}_{\alpha}})=\sigma(\mathscr{C}^{*}_{\alpha})=\left\{\frac{\Gamma(\alpha+1)\Gamma(z+\frac{1}{p})}{\Gamma(\alpha+z+\frac{1}{p})}\,:\,z\in\overline{\mathbb{C}_{+}}\right\}\cup\{0\},$
for $1\leq p<\infty,$ see Theorem 5.3 and [1, Theorem 7.6].
7. Universality and closed invariant subspaces for the semigroup $T(t)$
The Invariant Subspace Problem (ISP) is one of the most important open problems in Operator Theory on Hilbert Spaces. This problem is about finding if every linear bounded operator on an infinite-dimensional separable complex Hilbert space has a non-trivial closed invariant subspace.
In 1960, G. C. Rota [39, 40] presented the concept of an operator whose lattice of invariant subspaces possesses a sufficiently rich structure to model any Hilbert space operator, and demonstrated the existence of such operators. These are the universal operators.
Definition 7.1.
Let $H$ be an infinite-dimensional separable complex Hilbert space and let $\mathcal{B}(H)$ denote the algebra of bounded linear operators on $H$. An operator $U\in\mathcal{B}(H)$ is called universal if for every $T\in\mathcal{B}(H)$, there exist $\lambda\in\mathbb{C}$, $M\in\operatorname{Lat}U$ (the space of closed invariant subspaces for $U$), and an isomorphism $\Theta:H\to M$ such that $\Theta T=\lambda U\mid_{M}\Theta$. In other words, $T$ is similar to $\lambda U\mid_{M}$.
Now, the similarity $\Theta$ maps invariant subspaces of $T$ to those of $U\mid_{M}$. Suppose $U$ is a universal operator for a separable, infinite-dimensional complex Hilbert space $H$. Then every bounded operator on $H$ has an invariant subspace if and only if every infinite-dimensional subspace $M$ of $H$ that is invariant under $U$ has a non-zero proper subspace $M_{0}$ that is also invariant under $U$. Thus, we have the following known characterization.
Theorem 7.2.
Let $H$ be a separable Hilbert space and $U\in\mathcal{B}(H)$ be a universal operator. Then the following conditions are equivalent:
  1. (i)
    Every operator $T\in\mathcal{B}(H)$ has a non-trivial closed invariant subspace.
  2. (ii)
    Every minimal non-trivial closed invariant subspace of $U$ is one-dimensional.
Following the previous theorem, we say that a non-trivial closed invariant subspace for $U$ is minimal if it contains no proper non-trivial closed invariant subspaces for $U$. Consequently, understanding the Invariant Subspace Problem for Hilbert spaces reduces to understanding the invariant subspaces of the single operator $U,$ because all infinite-dimensional separable Hilbert spaces are isomorphic.
In 1969, S. R. Caradus [17] established a systematic tool to produce universal operators on Hilbert spaces.
Theorem 7.3.
Let $H$ be a separable Hilbert space and $T\in\mathcal{B}(H)$ such that:
  1. (i)
    $\dim\ker T=\infty$,
  2. (ii)
    and $T$ is surjective,
then $T$ is universal for $H$.
In the following we see that the properties of the spectrum of $T(t)$ studied in Section 4 give as a consequence the universality of translations of $T(t)$ on $L^{2}[0,1]$. Additionally, we will present several results concerning the minimal subspaces of $T(t)$. For convenience, we denote $\mathbb{S}:=\{\lambda\in\mathbb{C}\,:\,{\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/2\}.$
Theorem 7.4.
Let $t>0$ and $\mu\in B(0,e^{-t/2}),$ then the operator $\mu-T(t)$ is universal on $L^{2}[0,1]$. In particular, if every minimal non-trivial closed invariant subspace of $T(t)$ is one-dimensional for some $t>0$, then the Invariant Subspace Problem has a positive answer.
Proof.
On one hand, as we have seen in the proof of Theorem 4.1 we have that $\ker(\mu-T(t))$ is infinite-dimensional. In particular, if $\mu=e^{\lambda t}$ with $\lambda\in\mathbb{S},$
$\ker(\mu-T(t))=\overline{\operatorname{span}\{g_{\lambda+2k\pi i}\,:\,k\in\mathbb{Z}\}}^{L^{2}[0,1]}.$
On the other hand, as $\sigma_{r}(T(t))=\emptyset$ (see item (iv) of Theorem 4.1) we have that $\operatorname{Ran}(\mu-T(t))$ is dense. Lastly, by the Closed Range Theorem, (see [46, page 205] for a reference) $\operatorname{Ran}(\mu-T(t))$ is closed if and only $\operatorname{Ran}(\mu-S(t))$ is closed. Since $\sigma_{point}(S(t))=\emptyset$ and $\sigma_{ap}(S(t))=\delta B(0,e^{-t/2})$ (see Theorem 4.2) this means that for $\mu\in B(0,e^{-t/2})$, $\operatorname{Ran}(\mu-S(t))$ is closed and thus $\operatorname{Ran}(\mu-T(t))$ is closed as well. The application of the Caradus criterion stated before finishes the proof.
Lastly, as the closed invariant subspaces of $\mu-T(t)$ and $T(t)$ are the same, Theorem 7.2 finishes the proof, or merely taking $\mu=0.$
As discussed in the introduction, our goal is to understand the eigenfunctions and closed invariant subspaces of the operator $T(t)$ for $t>0$. The subspaces
$M_{r}:=M_{2,r}=\{f\in L^{2}[0,1]\mid f(s)=0\text{ a.e. for }s\in(0,r)\},\quad r\in(0,1),$
defined at the beginning of Section 4, are invariant under $T(t)$. Functions in $M_{e^{-t}}$ are all eigenfunctions with eigenvalue 0. Moreover, for each $r\in(0,1)$, $M_{r}\cap M_{e^{-t}}\neq\emptyset$, implying that every closed invariant subspace $M_{r}$ contains an eigenfunction. In particular, $T(t)$ is nilpotent on every closed invariant subspace $M_{r}$ with $r\in(0,1).$
Furthermore, for $\lambda\in\mathbb{S}$, the functions of the form $g_{\lambda+2k\pi i},$ $k\in\mathbb{Z},$ are eigenfunctions corresponding to the eigenvalue $e^{t\lambda}\in B(0,e^{-t/2})$. Consequently, the subspace spanned by each such eigenfunction is also invariant under $T(t)$. In particular, from the full Müntz–Szász theorem in $L^{2}[0,1]$ (see [42]) one can deduce that if $\{s_{n}\}_{n\in\mathbb{N}}\subset\mathbb{S},$ then the closed invariant subspaces
$\overline{\operatorname{span}\{g_{\lambda}\,:\,\lambda\in\{s_{n}\}_{n\in\mathbb{N}}\}}^{L^{2}[0,1]}$
of $T(t)$ are non-trivial if and only if
$\sum_{n=0}^{\infty}\frac{-(2{\mathfrak{R}}{\mathfrak{e}}\,s_{n}+1)}{|s_{n}|^{2}}<\infty.$
In [3, Proposition 3.8] the authors also showed that for $\lambda\in\mathbb{S},$ the $(m+1)^{\text{st}}$ generalized eigenspace of $\mathcal{C}$ ($\mathcal{C}_{1}$) with eigenvalue $\frac{1}{\lambda}$ lies in the linear span of $\{g_{\lambda},g_{\lambda}\log,\ldots,g_{\lambda}\log^{m}\},$ that is,
$\ker(\mathcal{C}+\frac{1}{\lambda})^{m+1}=\operatorname{span}\{g_{\lambda},g_{\lambda}\log,\ldots,g_{\lambda}\log^{m}\}.$
Similarly, for $t>0$ given, one can get for $\lambda\in\mathbb{S}$ that the $(m+1)^{\text{st}}$ generalized eigenspace of $T(t)$ with eigenvalue $e^{\lambda t}$ is the nontrivial closed invariant subspace
$\ker(e^{\lambda t}-T(t))^{m+1}=\overline{\operatorname{span}\{g_{\lambda+2k\pi i},g_{\lambda+2k\pi i}\log,\ldots,g_{\lambda+2k\pi i}\log^{m}\,:\,k\in\mathbb{Z}\}}^{L^{2}[0,1]},$
since in general $T(t)g_{\lambda}\log^{m}=e^{\lambda t}g_{\lambda}\sum_{j=0}^{m}(-t)^{m-j}\log^{j}.$
The natural next question is whether there are closed invariant subspaces of $T(t)$ beyond those already identified. Notably, the closed invariant subspaces identified so far are independent of $t$, prompting us to ask if there are any that depend on the value of $t$.
We now turn our attention to the relationship between the closed invariant subspaces of $T(t)$ and the Cesàro-Hardy operator defined on $L^{2}[0,1]$ by
$\mathcal{C}f(s)=\frac{1}{s}\int_{0}^{s}f(u)\,du=\int_{0}^{\infty}T(t)f(s)\,dt,\quad s\in(0,1].$
Note that for a closed subspace $M$ invariant under $T(t)$ for all $t>0$, $M$ is also invariant under $\mathcal{C}$. In other words, this means that
(7.1) $\cap_{t>0}\operatorname{Lat}T(t)\subseteq\operatorname{Lat}\mathcal{C}.$
The closed invariant subspaces of $\mathcal{C}$ have been recently characterized in [3, Theorem 1.4], introducing the concept of monomial space, which we present in the following lines.
Recall that $\lambda\in\mathbb{S}$ if and only if $g_{\lambda}\in L^{2}[0,1]$. For $S$ a finite subset of $\mathbb{S}$ we let $\mathcal{M}(S)$ denote the span in $L^{2}[0,1]$ of the monomials $g_{\lambda}$ whose exponents lie in $S$, i.e.,
$\mathcal{M}(S)=\left\{\sum_{\lambda\in S}a(\lambda)g_{\lambda}\,:\,a(\lambda)\in\mathbb{C}\text{ for }\lambda\in S\right\}.$
We refer to sets of $L^{2}[0,1]$ of this form as finite monomial spaces. The following definition appears in [3].
Definition 7.5.
We say that a subspace $\mathcal{M}$ of $L^{2}[0,1]$ is a monomial space if there exists a sequence $\mathcal{M}_{n}$ of finite monomial spaces such that $\mathcal{M}_{n}\to\mathcal{M}$ as $n\to\infty$, i.e.,
$\mathcal{M}=\left\{f\in L^{2}[0,1]\,:\,\lim_{n\to\infty}\operatorname{dist}(f,\mathcal{M}_{n})=0\right\}.$
Equipped with these definitions we can state [3, Theorem 1.4], adding the characterization of the closed subspaces which are invariant under $T(t)$ for every $t>0$.
Theorem 7.6.
Let $M$ be a closed non-zero subspace of $L^{2}[0,1]$. Then the following are equivalent:
  • (i)
    $M\in\operatorname{Lat}\mathcal{C}$,
  • (ii)
    $M$ is a monomial space,
  • (iii)
    $M\in\cap_{t>0}\operatorname{Lat}T(t)$.
Proof.
The equivalence between items (i) and (ii) is established in [3, Theorem 1.4]. The implication (iii) $\implies$ (i) follows from (7.1).
To show (ii) $\implies$ (iii), let $M$ be a monomial space in $L^{2}[0,1]$ and $f\in M$. By definition, there exists a sequence of finite monomial spaces $\mathcal{M}_{n}$ such that $\mathcal{M}_{n}\to\mathcal{M}$. Since $T(t)\mathcal{M}_{n}\subseteq\mathcal{M}_{n}$ (as $\mathcal{M}_{n}$ is the span of eigenfunctions of $T(t)$), we have:
$\operatorname{dist}(T(t)f,\mathcal{M}_{n})$ $=\inf_{m\in\mathcal{M}_{n}}\left\|T(t)f-m\right\|\leq\inf_{m^{\prime}\in T(t)\mathcal{M}_{n}}\left\|T(t)f-m^{\prime}\right\|$
$=\inf_{m\in\mathcal{M}_{n}}\left\|T(t)f-T(t)m\right\|\leq\left\|T(t)\right\|\inf_{m\in\mathcal{M}_{n}}\left\|f-m\right\|\to 0.$
Thus, $T(t)f\in M$ for every $t>0$, completing the proof. ∎
Remark $7.$
It is known that for a contraction $C_{0}$-semigroup, the lattice of the closed invariant subspaces of the semigroup is the same that the lattice of the closed invariant subspaces of the cogenerator (see [25, Chap. 2, Theorem 10.9]). In particular, $(e^{t}T(2t))_{t\geq 0}$ is contraction $C_{0}$-semigroup on $L^{2}[0,1]$ generated by $2A+I,$ and whose cogenerator is $V=((2A+I)+I)((2A+I)-I)^{-1},$ see Proposition 2.1 and Theorem $2.$ Therefore, by Remark 5.5 one can write
$2\mathcal{C}=4((2A+I)-I)^{-1}=V-I,$
and the equivalence between items (i) and (iii) on Theorem 7.6 can be also deduced from this remark.
To conclude this section, we present results on the minimal invariant subspaces of $T(t)$ for a fixed $t>0$. To know the minimal invariant subspaces of $T(t)$ would help to come near to the solution of the ISP. It is known that if $M$ is a minimal invariant subspace of $T(t)$, then
$M=K_{f}:=\overline{\operatorname{span}\{f,T(t)f,T(t)^{2}f,\dots\}}^{L^{2}[0,1]}$
for some nonzero $f\in M$. Therefore, to study minimal invariant subspaces of $T(t)$, we need to understand the cyclic subspaces $K_{f}$ for $f\in L^{2}[0,1]$.
Theorem 7.8.
Let $\lambda\in\mathbb{C}$ such that ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/2,$ $g=g_{\lambda}f$ for some $f\in L^{2}[0,1],$ and $L\in\mathbb{C}\setminus\{0\}$ such that $\lim_{s\to 0^{+}}f(s)=L$. Then $g_{\lambda}\in K_{g}$. In particular $K_{g}$ is minimal if and only if $f$ is a non-zero constant, that is, $K_{g}$ is one dimensional generated by the eigenfunction $g_{\lambda}.$
Proof.
First note that $g_{\lambda}f\in L^{2}[0,1]$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/2$ since $\lim_{s\to 0^{+}}f(s)$ is a complex number.
Now consider the sequence $(e^{-nt\lambda}(T(t)^{n}g)_{n\in\mathbb{N}},$ which is clearly contained in $K_{g}$ and fix $\varepsilon>0$. Notice that,
$\int_{0}^{1}\left|e^{-nt\lambda}T(nt)g(s)-Lg_{\lambda}(s)\right|^{2}\,ds$ $=\int_{0}^{1}\left|e^{-nt(\lambda+1)}g(e^{-nt}s)-Lg_{\lambda}(s)\right|^{2}\,ds$
$=\int_{0}^{1}\left|g_{\lambda}(s)\right|^{2}\left|f(e^{-nt}s)-L\right|^{2}\,ds.$
Now, as $\lim_{s\to 0^{+}}f(s)=L\neq 0$, there exists $\delta>0$ such that if $u\in(0,\delta)$ then $\left|f(u)-L\right|<\sqrt{-\varepsilon(2{\mathfrak{R}}{\mathfrak{e}}\,\lambda+1)}$. Take $n$ big enough such that $e^{-nt}<\delta$, then
$\int_{0}^{1}\left|e^{-nt\lambda}T(nt)g(s)-Lg_{\lambda}(s)\right|^{2}\,ds$ $\leq-\varepsilon(2{\mathfrak{R}}{\mathfrak{e}}\,\lambda+1)\int_{0}^{1}\left|g_{\lambda}(s)\right|^{2}\,ds=\varepsilon.$
So, $Lg_{\lambda}\in K_{g},$ and $K_{g}$ is minimal if and only if $f$ is a non-zero constant because the subspace generated by $g_{\lambda}$ is one-dimensional. ∎
Theorem 7.9.
Let $\lambda\in\mathbb{C}$ such that ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/2,$ $m\in\mathbb{N},$ $g=g_{\lambda}\log^{m}f$ for some $f\in L^{2}[0,1],$ and $L\in\mathbb{C}\setminus\{0\}$ such that $\lim_{s\to 0^{+}}f(s)=L$. Then $g_{\lambda}\in K_{g}.$ In particular $K_{g}$ contains the eigenfunction $g_{\lambda},$ and therefore it is not a minimal closed invariant subspace.
Proof.
First note that $g_{\lambda}\log^{m}f\in L^{2}[0,1]$ for ${\mathfrak{R}}{\mathfrak{e}}\,\lambda<-1/2$ since $\lim_{s\to 0^{+}}f(s)$ is a complex number.
Now consider the sequence $(\frac{e^{-nt\lambda}T(nt)g}{(nt)^{m}})_{n\in\mathbb{N}},$ which is contained in $K_{g}$ and fix $\varepsilon>0$. Denote by $h_{n}=\frac{e^{-nt\lambda}T(nt)(g_{\lambda}\log^{m})}{(nt)^{m}}.$ On one hand
$\int_{0}^{1}\left|\frac{e^{-nt\lambda}T(nt)g(s)}{(nt)^{m}}-Lh_{n}(s)\right|^{2}ds=\frac{1}{(nt)^{2m}}\int_{0}^{1}\left|g_{\lambda}(s)\log^{m}(e^{-nt}s)\right|^{2}\left|f(e^{-nt}s)-L\right|^{2}ds.$
As $\lim_{s\to 0^{+}}f(s)=L\neq 0$, there exists $\delta>0$ such that if $s\in(0,\delta)$ then $\left|f(s)-L\right|<\varepsilon$. Take $n$ big enough such that $e^{-nt}<\delta$, then
$\int_{0}^{1}\left|\frac{e^{-nt\lambda}T(nt)g(s)}{(nt)^{m}}-Lh_{n}(s)\right|^{2}\,ds$ $\leq\frac{\varepsilon^{2}}{(nt)^{2m}}\int_{0}^{1}\left|g_{\lambda}(s)\log^{m}(e^{-nt}s)\right|^{2}\,ds$
$=\varepsilon^{2}\int_{0}^{1}\left|g_{\lambda}(s)\right|^{2}\left|\sum_{j=0}^{m}\frac{(-1)^{m-j}\log^{j}(s)}{(nt)^{j}}\right|^{2}$
$\leq\varepsilon^{2}\sum_{j=0}^{m}\frac{1}{t^{j}}\int_{0}^{1}\left|g_{\lambda}(s)\log^{j}(s)\right|^{2}\,ds$
$=C\varepsilon^{2},$
with $C=\sum_{j=0}^{m}\frac{1}{t^{j}}\int_{0}^{1}\left|g_{\lambda}(s)\log^{j}(s)\right|^{2}\,ds.$
On the other hand,
$\int_{0}^{1}\left|Lh_{n}(s)-(-1)^{m}Lg_{\lambda}(s)\right|^{2}\,ds$ $=L^{2}\int_{0}^{1}\left|g_{\lambda}(s)\right|^{2}\left|\frac{\log^{m}(e^{-nt}s)}{(nt)^{m}}-(-1)^{m}\right|^{2}\,ds$
$=L^{2}\int_{0}^{1}\left|g_{\lambda}(s)\right|^{2}\left|\sum_{j=1}^{m}\frac{(-1)^{m-j}\log^{j}(s)}{(nt)^{j}}\right|^{2}\,ds$
$\leq L^{2}\sum_{j=1}^{m}\frac{1}{(nt)^{j}}\int_{0}^{1}\left|g_{\lambda}(s)\log^{j}(s)\right|^{2}\,ds\to 0,$
as $n\to\infty.$ So we conclude that $Lg_{\lambda}\in K_{g}.$